# ED-ViT: Splitting Vision Transformer for Distributed Inference on Edge Devices

### v1

@@
Abstract.
Deep learning models are increasingly deployed on resource-constrained edge devices for real-time data analytics. In recent years, Vision Transformer models and their variants have demonstrated outstanding performance across various computer vision tasks. However, their high computational demands and inference latency pose significant challenges for model deployment on resource-constraint edge devices. To address this issue, we propose a novel Vision Transformer splitting framework, ED-ViT, designed to execute complex models across multiple edge devices efficiently. Specifically, we partition Vision Transformer models into several sub-models, where each sub-model is tailored to handle a specific subset of data classes. To further minimize computation overhead and inference latency, we introduce a class-wise pruning technique that reduces the size of each sub-model. We conduct extensive experiments on five datasets with three model structures, demonstrating that our approach significantly reduces inference latency on edge devices and achieves a model size reduction of up to 28.9 times and 34.1 times, respectively, while maintaining test accuracy comparable to the original Vision Transformer. Additionally, we compare ED-ViT with two state-of-the-art methods that deploy CNN and SNN models on edge devices, evaluating accuracy, inference time, and overall model size. Our comprehensive evaluation underscores the effectiveness of the proposed ED-ViT framework.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computer systems organization Embedded softwareccs: Computing methodologies Computer vision tasksccs: Computing methodologies Neural networks
1. Introduction
@@
In recent years, deep learning models have been increasingly deployed on resource-constrained edge devices to meet the growing demand for real-time data analytics in industrial systems  (Chen et al., 2021, 2023; Yu et al., 2024) and have demonstrated remarkable capabilities in various applications such as video image analysis and speech recognition. Convolutional Neural Networks (CNNs) (Krizhevsky et al., 2012) like VGGNet (Simonyan and Zisserman, 2014) and ResNet (He et al., 2015), as well as Spike Neural Networks (SNNs) (Deng and Gu, 2021), have achieved satisfactory performance in many edge computing scenarios. As the field progresses, researchers are exploring the deployment of more complex structured models on edge devices to further improve performance. Transformer architecture (Vaswani, 2017), which has revolutionized natural language processing (NLP) tasks, has inspired similar advancements in computer vision. Vision Transformer (ViT) models (Alexey, 2020) and their variants have shown outstanding results across various computer vision tasks, including image classification (Wu et al., 2020; Touvron et al., 2021), object detection (Carion et al., 2020; Dai et al., 2021a; Yang et al., 2021b), semantic segmentation (Song et al., 2020; Wang et al., 2021c) and action recognition (Girdhar et al., 2019; Plizzari et al., 2021) and audio spectrogram recognition (Gong et al., 2022). The success of ViTs has sparked interest in leveraging their capabilities for edge computing applications.
@@
However, the rapid advancement in machine learning technologies has increased the demand for computational resources and memory, given the complexity of these model configurations. Achieving higher accuracy with ViT requires substantial computational power and memory, which poses challenges for deployment on edge devices. For instance, ViT-Base (Alexey, 2020) consists of over 86.7 million parameters and requires approximately 330MB of memory. Researchers now face the dilemma of deploying such complex models while dealing with resource-constrained resources.
@@
Previous studies aiming to reduce the deployment overhead primarily focus on compressing Vision Transformer models. These approaches can be classified into three major categories: (1) architecture and hierarchy restructuring (Pan et al., 2021; Graham et al., 2021), (2) encoder block enhancements (Zhang et al., 2021; Yu et al., 2021; Yang et al., 2021a; Tu et al., 2022; Yao et al., 2022; Pan et al., 2023), and (3) integrated approaches (Yuan et al., 2021; Dai et al., 2021b). However, these methods often suffer from either poor inference accuracy or high inference latency as they attempt to compress a large model to fit into a memory-constrained edge device.
@@
Figure 1. The overview of ED-ViT which consists of four steps: Model Splitting, Model Pruning, Model Assignment and Model Fusion.
@@
To develop a solution that mitigates accuracy drop and enables efficient deployment of the Vision Transformer on resource-constraint edge devices, we aim to utilize the collaboration of multiple edge devices and propose a Vision Transformer splitting framework Edge Device Vision Transformer called ED-ViT. As illustrated in Figure 1, ED-ViT first partitions the original Vision Transformer into several smaller sub-models, inspired by the concept of split learning (SL). However, unlike traditional SL, which does not consider edge device constraints, each of these small sub-models is responsible for detecting a specific subset of the classes and is deployed on resource-constrained edge devices. ED-ViT then employs model pruning techniques to further alleviate the computational load and processing requirement for each sub-model. To optimize model assignment, we design a greedy assignment algorithm that takes into account both the model computational resources and memory resources. Besides integrating the previous steps, ED-ViT uses a multilayer perceptron (MLP) model to fuse the results from all the sub-models. We conduct experiments across five datasets to evaluate the effectiveness of ED-ViT framework on edge devices, particularly in low-power video analytics. The results, measured across three key metrics—accuracy, inference latency, and model size—consistently demonstrate the significant benefits of ED-ViT. Additionally, we compare ED-ViT with methods that split CNN and SNN, highlighting the great potential of deploying Vision Transformer on resource-constrained edge devices to achieve high accuracy while maintaining small model sizes and low inference latency. Our main contributions are summarized as follows:
@@
  • This is the first study focused on splitting and deploying Vision Transformer onto edge devices. We propose a framework that leverages the capabilities of Vision Transformer, allowing for the collaboration of multiple edge devices to achieve distributed inference in practical applications.
@@
  • We conduct extensive experiments with three computer vision datasets and two audio recognition datasets across three ViT structures, demonstrating that our framework significantly reduces inference latency on edge devices and decreases overall memory usage with negligible accuracy loss in various Vision Transformer applications.
2. Related Works
2.1. General Vision Transformer Compression
@@
Local pruning techniques focus on removing redundant components within specific layers of the model. For instance, PVT (Wang et al., 2021b) and its successor PVTv2 (Wang et al., 2021a) introduce a pyramid hierarchical structure to transformer backbones, achieving high accuracy with reduced computation. Zhu et al. (2021) applies sparsity regularization during training and subsequently prunes the dimensions of linear projections, targeting less significant parameters. Xia et al. (2022) prunes multi-head self-attention (MHSA) and feed-forward networks (FFN), which are often redundant components.  Liang et al. (2022); Liu et al. (2023) proposes network pruning to eliminate complexity and model sizes by reducing tokens. Other noteworthy contributions include DToP (Tang et al., 2023), which enables early token exits for semantic segmentation tasks. Conversely, global pruning techniques adopt a comprehensive perspective by evaluating and pruning the overall significance of neurons or layers across the entire network. SAViT (Zheng et al., 2022) purposes structure-aware Vision Transformer pruning via collaborative optimization. For instance, CP-ViT (Song et al., 2022) systematically assesses the importance of head and attention layers for the purpose of pruning, while Evo-ViT (Xu et al., 2022) identifies and preserves significant tokens, thereby discarding those of lesser importance. Moreover, the Skip-attention approach (Venkataramanan et al., 2023) facilitates the omission of entire self-attention layers, thereby exemplifying a global pruning methodology. X-pruner (Yu and Xiang, 2023) employs explainability-aware masks to inform its pruning decisions, thereby advancing a more informed global pruning strategy. In addition, UP-ViT (Yu and Wu, 2023) introduces a unified pruning framework that leverages KL divergence to guide the decision-making process for pruning, while LORS (Li et al., 2024) optimizes parameter usage by sharing the majority of parameters across stacked modules, thereby necessitating fewer unique parameters per module.
Among existing pruning methods, UP-ViT (Yu and Wu, 2023) has the closest resemblance to our approach. However, it is important to note that these techniques cannot be directly applied to edge devices: they often suffer from poor performance when the pruning ratio is high or incur high computation overhead when the pruning ratio is low, making them unsuitable for resource-constrained edge environments. In contrast, our work introduces a class-based global structured pruning method that addresses these limitations. Our approach is orthogonal to most previous methods and does not involve trainable parameters, contributing to more stable performance.
2.2. Vision Transformer on Edge Devices
Architecture and Hierarchy Restructuring: HVT (Pan et al., 2021) compresses sequential resolutions using hierarchical pooling, reducing computational cost and enhancing model scalability. LeViT (Graham et al., 2021) is a hybrid model that combines the strengths of CNNs and transformers. For image classification tasks, it utilizes the hierarchical structure of LeNet (Krizhevsky et al., 2012) to optimize the balance between accuracy and efficiency, and uses average pooling in the feature map stage.
Encoder Block Enhancements: ViL (Zhang et al., 2021) introduces a multiscale vision longformer that lessens computational and memory complexity when encoding high-resolution images. Poolformer (Yu et al., 2021) deliberately replaces the attention module in transformers with a simple pooling layer. LiteViT (Yang et al., 2021a) introduces a compact transformer backbone with two new lightweight self-attention modules (self-attention and recursive atrous self-attention) to mitigate performance loss. Dual-ViT (Yao et al., 2022) reduces feature map resolution, consisting of two dual-block and two merge-block stages. MaxViT (Tu et al., 2022) divides attention into local and global components and decomposes it into a sparse form with window and grid attention. Slide-Transformer (Pan et al., 2023) proposes a slide attention module to address the problem that computational complexity increases quadratically with the attention modules.
Integrated Approaches: Some methods integrate both of the above approaches. CeiT (Yuan et al., 2021) combines Transformer and CNN strengths to overcome the shortcomings of each, incorporating an image-to-tokens module, locally-enhanced feedforward layers, and layer-wise class token attention. CoAtNet (Dai et al., 2021b) combines depth-wise convolutions and simplifies traditional self-attention by relative attention, enhancing efficiency by stacking convolutions and attention layers.
However, they fail to link pruning with specific classes, which limits their methods when both high performance and low memory usage are required.
2.3. Split Learning
Current works that combine Vision Transformer and split learning primarily focus on federated learning, addressing data privacy and efficient collaboration in multi-client environments (Oh et al., 2022; Almalik et al., 2023; Oh et al., 2024), where the inner structure of a large model is split across smaller devices and later fused (Su et al., 2022). However, these approaches do not target the deployment of Vision Transformer on edge devices.
Traditional machine learning model splitting generally involves partitioning a large model into multiple smaller sub-models that can be executed collaboratively on resource-constrained devices, providing a promising technique for deploying models on edge devices. Splitnet (Kim et al., 2017) clusters classes into groups, partitioning a deep neural network into tree-structed sub-networks. Bakhtiarnia et al. (2022) dynamically partitions models based on the communication channel’s state. Nnfacet (Chen et al., 2021, 2023) splits large CNNs into lightweight class-specific sub-models to accommodate device memory and energy constraints, with the sub-models being fused later. Yu et al. (2024) follows a similar approach to split deep SNNs across edge devices. Distredge (Hou et al., 2022) uses deep reinforcement learning to compute the optimal partition for CNN models.
@@ 2.3. Split Learning
Model depth width heads params Flops Lantency Mem Size
@@
Table 1. Performance characteristics for standard Vision Transformer models with their default parametrization at resolution 224 $\times$ 224: ViT-Small, ViT-Base and ViT-Large, all with patch size of 16 $\times$ 16. All the experiment results are obtained on Raspberry Pi-4B.
@@
To the best of our knowledge, our work presents the first exploration of Vision Transformer model partitioning for edge deployment, marking a significant contribution to this field. Drawing inspiration from previous studies (Chen et al., 2021, 2023; Yu et al., 2024), our framework, ED-ViT, introduces an innovative approach to decompose a multi-class ViT model into several class-specific sub-models, each performing a subset of classification. Unlike relying on channel-wise pruning, ED-ViT employs advanced pruning techniques specifically designed for the unique architecture of Vision Transformers.
3. Problem Formulation
@@
The structures of three representative Vision Transformer models, ViT-Small, ViT-Base, and ViT-Large, are presented in Table 1. The number of operations is commonly used to estimate computational energy consumption at the hardware level. In Vision Transformer models, almost all floating-point operations (FLOPs) are multiply-accumulate (MAC) operations.
@@
Thus, we formulate the problem as follows. We assume that we have $L$ inference samples in total to be processed, and the set of $N$ edge devices is represented as $D$. The available memory and energy (FLOPs of an edge device) (Weaver, 2024) for each edge device $D_{i}$ are denoted as $M_{i}$ and $E_{i}$, respectively. The FLOPs (energy consumption) for each inference sample for the sub-model $Model_{j}$ from the set of sub-models $\{Models\}$ is represented as $e_{j}$, calculated based on the previous energy consumption estimation. To formulate the problem of Vision Transformer partitioning and edge-device-based deployment, we define the objective function as $max_{\{Model_{j}\}}min_{D_{i}\in D}\{E_{i}-Le_{j}\}$, aiming to minimize the maximal inference latency, as inference latency is closely related to the computational power of edge devices. Additionally, the accuracy $a_{fus}$ of the fused results from all $N$ inference samples must be greater than or equal to the required inference accuracy $A_{re}$; the total memory sizes of all sub-models should not exceed the memory budget $bu$.
@@
The optimization problem can be formally formulated as follows, where $x_{ie}$ is a binary decision variable: 1 indicates that the sub-model deployed on edge device $D_{i}$ is responsible for class $e$, and 0 otherwise. Each sub-model learns a specific subset of the classes in $C$. Furthermore, the memory consumption of sub-model $j$, denoted as $size$(Modelj), represented as $m_{j}$, must be smaller than the available memory size of the deployed edge device:
@@
(1) {$&max_{\{Model_{j}\}}min_{D_{i}\in D}\{E_{i}-Le_{j}\}\\ &\text{s.t.}\quad Le_{j}\leq E_{i},\quad\text{Model}_{j}\quad\text{deploys on}\quad D_{i}\\ &\quad\quad m_{j}\leq M_{i},\quad\\ &\quad\quad a_{fus}\geq A_{re},\\ &\quad\quad\sum_{j}m_{j}\leq bu,\\ &\quad\quad\sum_{i=1}^{|D|}x_{ie}=1,\quad\forall e\in C,\forall i\in D\end{aligned}\right.$
@@
Figure 2. Structured pruning of a Vision Transformer block. Left: illustration of prunable components in a ViT block. Right: corresponding sequential pruning process. Our approach targets three key components: (1) channels in residual connections (red, denoted as $d$), (2) the number of heads in the MHSA module (green, denoted as $h$), and (3) hidden layer channels in the FFN (blue, denoted as $c$). The pruning process occurs in three stages: residual connection channels, MHSA heads, and FFN hidden dimensions. Yellow regions indicate parameters being pruned in the current stage, while gray regions represent previously pruned parameters.
4. Methodology

### v2

# Zemin Liu4, Linshan Jiang6, Jialin Li1

@@
Abstract
Deep learning models are increasingly utilized on resource-constrained edge devices for real-time data analytics. Recently, Vision Transformer and their variants have shown exceptional performance in various computer vision tasks. However, their substantial computational requirements and low inference latency create significant challenges for deploying such models on resource-constrained edge devices. To address this issue, we propose a novel framework, ED-ViT, which is designed to efficiently split and execute complex Vision Transformers across multiple edge devices. Our approach involves partitioning Vision Transformer models into several sub-models, while each dedicated to handling a specific subset of data classes. To further reduce computational overhead and inference latency, we introduce a class-wise pruning technique that decreases the size of each sub-model. Through extensive experiments conducted on five datasets using three model architectures and actual implementation on edge devices, we demonstrate that our method significantly cuts down inference latency on edge devices and achieves a reduction in model size by up to 28.9 times and 34.1 times, respectively, while maintaining test accuracy comparable to the original Vision Transformer. Additionally, we compare ED-ViT with two state-of-the-art methods that deploy CNN and SNN models on edge devices, evaluating metrics such as accuracy, inference time, and overall model size. Our comprehensive evaluation underscores the effectiveness of the proposed ED-ViT framework.
Index Terms:
@@ Abstract
I Introduction
@@
In recent years, deep learning models have been increasingly deployed on resource-constrained edge devices to meet the growing demand for real-time data analytics in industrial systems [1, 2, 3] and have demonstrated remarkable capabilities in various applications such as video image analysis and speech recognition. Convolutional Neural Networks (CNNs) [4] like VGGNet [5] and ResNet [6], as well as Spike Neural Networks (SNNs) [7], have achieved satisfactory performance in many edge computing scenarios. As the field progresses, researchers are exploring the deployment of more complex structured models on edge devices to further improve performance. Transformer architecture [8], which has revolutionized natural language processing (NLP) tasks, has inspired similar advancements in computer vision. Vision Transformer (ViT) models [9] and their variants have shown outstanding results across various computer vision tasks, including image classification [10, 11], object detection [12, 13, 14], semantic segmentation [15, 16] and action recognition [17, 18] and audio spectrogram recognition [19]. The success of ViTs has sparked interest in leveraging their capabilities for edge computing applications.
@@
However, the rapid advancement in machine learning technologies has increased the demand for computational resources and memory, given the complexity of these model configurations. Achieving higher accuracy with ViT requires substantial computational power and memory, which poses challenges for deployment on edge devices. For instance, ViT-Base [9] consists of over 86.7 million parameters and requires approximately 330MB of memory. Researchers now face the dilemma of deploying such complex models while dealing with resource-constrained resources.
@@
Previous studies aiming to reduce the deployment overhead primarily focus on compressing Vision Transformer models. These approaches can be classified into three major categories: (1) architecture and hierarchy restructuring [20, 21], (2) encoder block enhancements [22, 23, 24, 25, 26, 27], and (3) integrated approaches [28, 29]. However, these methods often suffer from either poor inference accuracy or high inference latency as they attempt to compress a large model to fit into a memory-constrained edge device.
@@
Figure 1: The overview of ED-ViT, including four steps: Model Splitting, Model Pruning, Model Assignment and Model Fusion.
@@
To develop a solution that mitigates accuracy drop and enables efficient deployment of the Vision Transformer on resource-constraint edge devices, we aim to utilize the collaboration of multiple edge devices and propose a Vision Transformer splitting framework named Edge Device Vision Transformer, shorten as ED-ViT. As illustrated in Fig. 1, ED-ViT first partitions the original Vision Transformer into several smaller sub-models, inspired by the concept of split learning (SL). However, unlike traditional SL, which does not consider edge device constraints, each of these small sub-models is responsible for detecting a specific subset of the classes and is deployed on resource-constrained edge devices. ED-ViT then employs model pruning techniques to further alleviate the computational load and processing requirement for each sub-model. To optimize model assignment, we design a greedy assignment algorithm that takes into account both the model computational resources and memory resources. Besides integrating the previous steps, ED-ViT uses a multilayer perceptron (MLP) model to fuse the results from all the sub-models. We conduct experiments across five datasets to evaluate the effectiveness of ED-ViT framework on edge devices, particularly in low-power video analytics. The results, measured across three key metrics—accuracy, inference latency, and model size—consistently demonstrate the significant benefits of ED-ViT. Additionally, we compare ED-ViT with methods that split CNN and SNN, highlighting the great potential of deploying Vision Transformer on resource-constrained edge devices to achieve high accuracy while maintaining small model sizes and low inference latency. Our main contributions are summarized as follows:
@@
  • This is the first study focusing on combining pruning and splitting, deploying Vision Transformer onto edge devices. We propose a framework that leverages the capabilities of Vision Transformer, allowing for the collaboration of multiple edge devices to achieve distributed inference in practical applications.
@@
  • We conduct extensive experiments with three computer vision datasets and two audio recognition datasets across three ViT structures. Besides, we implement our ED-ViT on Raspberry Pi 4Bs, demonstrating that our framework significantly reduces inference latency on edge devices and decreases overall memory usage with negligible accuracy loss in various applications.
II Related Works
II-A General Vision Transformer Compression
@@
TABLE I: Performance characteristics for standard Vision Transformer models with their default parametrization at resolution 224 $\times$ 224: ViT-Small, ViT-Base and ViT-Large, all with patch size of 16 $\times$ 16. All the experiment results are obtained on Raspberry Pi-4B devices.
@@
Model Depth Width Heads Params Flops Latency Mem Size
@@
Local pruning techniques focus on removing redundant components within specific layers of the model. For instance, PVT [30] and its successor PVTv2 [31] introduce a pyramid hierarchical structure to transformer backbones, achieving high accuracy with reduced computation. [32] applies sparsity regularization during training and subsequently prunes the dimensions of linear projections, targeting less significant parameters. [33] prunes multi-head self-attention (MHSA) and feed-forward networks (FFN), which are often redundant components. [34, 35] propose network pruning to eliminate complexity and model sizes by reducing tokens. Other noteworthy contributions include DToP [36], which enables early token exits for semantic segmentation tasks. Conversely, global pruning techniques adopt a comprehensive perspective by evaluating and pruning the overall significance of neurons or layers across the entire network. SAViT [37] purposes structure-aware Vision Transformer pruning via collaborative optimization. For instance, CP-ViT [38] systematically assesses the importance of head and attention layers for the purpose of pruning, while Evo-ViT [39] identifies and preserves significant tokens, thereby discarding those of lesser importance. Moreover, the Skip-attention approach [40] facilitates the omission of entire self-attention layers, thereby exemplifying a global pruning methodology. X-pruner [41] employs explainability-aware masks to inform its pruning decisions, thereby advancing a more informed global pruning strategy. In addition, UP-ViT [42] introduces a unified pruning framework that leverages KL divergence to guide the decision-making process for pruning, while LORS [43] optimizes parameter usage by sharing the majority of parameters across stacked modules, thereby necessitating fewer unique parameters.
Among existing pruning methods, UP-ViT [42] has the closest resemblance to our approach. However, it is important to note that these techniques cannot be directly applied to edge devices: they often suffer from poor performance when the pruning ratio is high or incur high computation overhead when the pruning ratio is low, making them unsuitable for resource-constrained edge environments. In contrast, our work introduces a class-based global structured pruning method that addresses these limitations. Our approach is orthogonal to most previous methods and does not involve trainable parameters, contributing to more stable performance.
II-B Vision Transformer on Edge Devices
Architecture and Hierarchy Restructuring: HVT [20] compresses sequential resolutions using hierarchical pooling, reducing computational cost and enhancing model scalability. LeViT [21] is a hybrid model that combines the strengths of CNNs and transformers. For image classification tasks, it utilizes the hierarchical structure of LeNet [4] to optimize the balance between accuracy and efficiency, and uses average pooling in the feature map stage. MobileViTv3 [44] propose changes to the fusion block, which addresses the scaling and simplifies the learning tasks.
Encoder Block Enhancements: ViL [22] introduces a multiscale vision longformer that lessens computational and memory complexity when encoding high-resolution images. Poolformer [23] deliberately replaces the attention module in transformers with a simple pooling layer. LiteViT [24] introduces a compact transformer backbone with two new lightweight self-attention modules (self-attention and recursive atrous self-attention) to mitigate performance loss. Dual-ViT [26] reduces feature map resolution, consisting of two dual-block and two merge-block stages. MaxViT [25] divides attention into local and global components and decomposes it into a sparse form with window and grid attention. Slide-Transformer [27] proposes a slide attention module to address the problem that computational complexity increases quadratically with the attention modules, while EdgeViT [45] enables attention-based vision models to compete with the best light-weight CNNs when considering the tradeoff between accuracy and on-device efficiency.
Integrated Approaches: Some methods integrate both of the above approaches. CeiT [28] combines Transformer and CNN strengths to overcome the shortcomings of each, incorporating an image-to-tokens module, locally-enhanced feedforward layers, and layer-wise class token attention. CoAtNet [29] combines depth-wise convolutions and simplifies traditional self-attention by relative attention, enhancing efficiency by stacking convolutions and attention layers. DeViT [46] also decomposes Vision Transformer for collaborative inference. However, DeViT trains a ViT-Large for each sub-model even when splitting ViT-Small and employs model distillation to enhance accuracy, which introduces significant training overhead. In addition, the smallest model size that DeViT provides is larger than 90MB.
However, they never consider linking pruning with specific classes, which limits their methods when both high performance and low memory usage are required.
II-C Split Learning
Current works that combine Vision Transformer and split learning primarily focus on federated learning, addressing data privacy and efficient collaboration in multi-client environments [47, 48, 49], where the inner structure of a large model is split across smaller devices and later fused [50]. However, these approaches do not target the deployment of Vision Transformer on edge devices.
Traditional machine learning model splitting generally involves partitioning a large model into multiple smaller sub-models that can be executed collaboratively on resource-constrained devices, providing a promising technique for deploying models on edge devices. Splitnet [51] clusters classes into groups, partitioning a deep neural network into tree-structed sub-networks. [52] dynamically partitions models based on the communication channel’s state. Nnfacet [1, 2] splits large CNNs into lightweight class-specific sub-models to accommodate device memory and energy constraints, with the sub-models being fused later. [3] follows a similar approach to split deep SNNs across edge devices. Distredge [53] uses deep reinforcement learning to compute the optimal partition for CNN models.
@@ II-C Split Learning
To the best of our knowledge, our work presents the first exploration of Vision Transformer model partitioning for edge deployment, marking a significant contribution to this field. Drawing inspiration from previous studies [1, 2, 3], our framework, ED-ViT, introduces an innovative approach to decompose a multi-class ViT model into several class-specific sub-models, each performing a subset of classification. Unlike relying on channel-wise pruning, ED-ViT employs advanced pruning techniques specifically designed for the unique architecture of Vision Transformers.
III Problem Formulation
@@
The structures of three representative Vision Transformer, ViT-Small, ViT-Base, and ViT-Large, are presented in Table I. The number of operations is commonly used to estimate computational energy consumption at the hardware level. In Vision Transformer models, almost all floating-point operations (FLOPs) are multiply-accumulate (MAC) operations.
@@
Thus, we formulate the problem as follows. We assume that we have $L$ inference samples in total to be processed, and the set of $N$ edge devices is represented as $D$. The available memory and energy (FLOPs of an edge device) [54] for each edge device $D_{i}$ are denoted as $M_{i}$ and $E_{i}$, respectively. The FLOPs (energy consumption) for each inference sample for the sub-model $Model_{j}$ from the set of sub-models $\{Models\}$ is represented as $e_{j}$, calculated based on the previous energy consumption estimation. To formulate the problem of Vision Transformer partitioning and edge-device-based deployment, we define the objective function as $max_{\{Model_{j}\}}min_{D_{i}\in D}\{E_{i}-Le_{j}\}$, aiming to minimize the maximal inference latency, as inference latency is closely related to the computational power of edge devices. Additionally, the accuracy $a_{fus}$ of the fused results from all $N$ inference samples must be greater than or equal to the required inference accuracy $A_{re}$; the total memory sizes of all sub-models should not exceed the memory budget $bu$.
@@
The optimization problem can be formally formulated as follows, where $x_{ie}$ is a binary decision variable: $1$ indicates that the sub-model deployed on edge device $D_{i}$ is responsible for class $e$, and 00 otherwise. Each sub-model learns a specific subset of the classes in $C$. Furthermore, the memory consumption of sub-model $j$, denoted as $size$(Modelj), represented as $m_{j}$, must be smaller than the available memory size of the deployed edge device:
@@
{$&argmax_{\{Model_{j}\}}min_{D_{i}\in D}\{E_{i}-Le_{j}\}\\ &\text{s.t.}\quad Le_{j}\leq E_{i},\quad\text{Model}_{j}\quad\text{deploys on}\quad D_{i}\\ &\quad\quad m_{j}\leq M_{i},\quad\\ &\quad\quad a_{fus}\geq A_{re},\\ &\quad\quad\sum_{j}m_{j}\leq bu,\\ &\quad\quad\sum_{i=1}^{|D|}x_{ie}=1,\quad\forall e\in C,\forall i\in D\end{aligned}\right.$ (1)
@@
Figure 2: Structured pruning of a Vision Transformer block. Left: illustration of prunable components in a ViT block. Right: corresponding sequential pruning process. Our approach targets three key components: (1) channels in residual connections (red, denoted as $d$), (2) the number of heads in the MHSA module (green, denoted as $h$), and (3) hidden layer channels in the FFN (blue, denoted as $c$). The pruning process occurs in three stages: residual connection channels, MHSA heads, and FFN hidden dimensions. Yellow regions indicate parameters being pruned in the current stage, while gray regions represent previously pruned parameters.
IV Methodology